OpenStudy (anonymous):

interal (dx/sqrt(x^(2)+16)

OpenStudy (solomonzelman):

\(\large\color{slate}{\displaystyle\int\limits_{~}^{~}\frac{1}{x^2+16}~dx}\) ?

OpenStudy (solomonzelman):

correct interpretation ?

OpenStudy (kl0723):

\[\int\limits \frac{ dx }{ \sqrt{x^2+16} }\] like this

OpenStudy (solomonzelman):

\(\large\color{slate}{\displaystyle\int\limits_{~}^{~}\frac{1}{\sqrt{x^2+16}}~dx}\)

OpenStudy (anonymous):

Yeah

OpenStudy (noelgreco):

First, factor the 16 out of the radical.

OpenStudy (solomonzelman):

what ?

OpenStudy (kl0723):

nope

OpenStudy (xapproachesinfinity):

this is one of those trig integrals

OpenStudy (kl0723):

yeah is the form \[\sqrt{x^2+a^2}\]

OpenStudy (misty1212):

HI!!

OpenStudy (xapproachesinfinity):

so what did you do so far?

OpenStudy (misty1212):

try \(x=4\tan(\theta)\)

OpenStudy (misty1212):

the radical will go bye bye lickety split

OpenStudy (kl0723):

yeah is x=aTan(angle), so as @misty1212 says a=4 and x=4Tan(theta)

OpenStudy (xapproachesinfinity):

well i guess this is solved :)

OpenStudy (kl0723):

the person that asked this question left :P and we're still trying to help *sighs*

OpenStudy (noelgreco):

\[\int\limits_{}^{}\frac{ 1 }{ 4\sqrt{(\frac{ x }{ 4 }})^{2}+1 }\]

OpenStudy (kl0723):

@NoelGreco a direct sub method is the way to go

OpenStudy (misty1212):

for a simpler method, memorize \[\frac{d}{dx}[\sinh^{-1}(x)]=\frac{1}{\sqrt{1+x^2}}\]

OpenStudy (xapproachesinfinity):

eh i'm against memorizing hehe i was going to say that arcsin' = that stuff but better learn that sub

OpenStudy (noelgreco):

Now you have it in the form x= tan theta, d/d theta =sec^2 theta.

OpenStudy (solomonzelman):

you can prove that what mitsy posted as well, if you like very simply.

OpenStudy (solomonzelman):

I do like memorizing:D

OpenStudy (solomonzelman):

why did jacket leave?

OpenStudy (solomonzelman):

this user is probably south from the equator jk

OpenStudy (xapproachesinfinity):

hehe good analysis :)

OpenStudy (xapproachesinfinity):

@SolomonZelman how to use show that one i haven't seen that much of arcsinh it just kind of remind me of arctan just the square root difference

OpenStudy (xapproachesinfinity):

how do you*

OpenStudy (solomonzelman):

you want arctan(s) or what?

OpenStudy (solomonzelman):

arctan(x) * ?

OpenStudy (solomonzelman):

\(\large\color{slate}{ \displaystyle \frac{d}{dx}\tan^{-1}x=1/( x^2+1) }\) ~~~~~~~~~~~~~~~~~~~~~~~~~ \(\large\color{slate}{ \displaystyle y=\tan^{-1}x }\) \(\large\color{slate}{ \displaystyle \tan y=x }\) derivative, \(\large\color{slate}{ \displaystyle y'~\sec^2y=1 }\) \(\large\color{slate}{ \displaystyle y'~(\tan^2y+1)=1 }\) \(\large\color{slate}{ \displaystyle y'~( x^2+1)=1 }\) \(\large\color{slate}{ \displaystyle y'=1/( x^2+1) }\)

OpenStudy (solomonzelman):

this simple thing you want?

OpenStudy (solomonzelman):

who is here? was it the world's biggest lag? jk

OpenStudy (solomonzelman):

igtg in a couple minutes it was nice to see the problem.

OpenStudy (xapproachesinfinity):

no i meant arcsinh i know arctan of course :)

OpenStudy (xapproachesinfinity):

i thought you were talking about arcsinh'

OpenStudy (xapproachesinfinity):

@SolomonZelman